OpenStudy (anonymous):

How do I solve this please help! 1/2y+8=1/6y

OpenStudy (anonymous):

times both sides by six

OpenStudy (anonymous):

multiply everything by 6: 3y + 48 = y subtract y from both sides: 2y + 48 = 0 subtract 48 from both sides: 2y=-48 divide by 2: y=-24

OpenStudy (anonymous):

\[y=\frac{-1}{24}\]

OpenStudy (anonymous):

........

OpenStudy (anonymous):

1/(2 y)+8 = 1/(6 y) Subtract 1/(6 y) from both sides: 1/(3 y)+8 = 0 Subtract 8 from both sides: 1/(3 y) = -8 Divide both sides by 1/3: 1/y = -24 Raise both sides to the power of -1: y = -1/24

OpenStudy (anonymous):

you have to get y in to one side, go on simply.

OpenStudy (anonymous):

i dont believe y is in the denominator dhashni

OpenStudy (anonymous):

dhashni you're making it way too complicated

OpenStudy (anonymous):

Are you using PEMDAS?You will have to simplfly

OpenStudy (anonymous):

no dhashni believes y is in the denominator\[\frac{1}{2y}+8=\frac{1}{6y}\]

OpenStudy (anonymous):

ally made it more easy for understand to desybaybee

OpenStudy (anonymous):

Ally I love your example!! But how is the 24 negative? I don't understand I did it and got positive???

OpenStudy (anonymous):

ally is correct if really what you have is \[\frac{1}{2}y+8=\frac{1}{6}y\]

OpenStudy (anonymous):

because you have 2y=-48 -48/2 = -24

OpenStudy (anonymous):

if you ever suspect your answer to be wrong or aren't 100% sure put it back into the equation and see if they equal each other

OpenStudy (anonymous):

negative divided by positive equal negative

OpenStudy (anonymous):

Ahhhh ! But which number was negative to begin with?

OpenStudy (anonymous):

I really hate math and I'm taking this pre course in college :(

OpenStudy (anonymous):

1/2y+8=1/6y okay so when you multiply by 6 you get: 3y + 48 = y then subtract y from both sides 3y - y + 48 = y - y 2y +48 = 0 then subtract 48 from both sides: 2y +48 -48 = 0 -48 2y = -48 divide both sides by 2 2y/2 = -28/2 y=-24

OpenStudy (anonymous):

Ohhhh LOL I feel stupid I skipped the step from subtracting 48 from both sides and went straight to dividing 2y=48 and thats why I was getting a positive answer

OpenStudy (anonymous):

So if I have the problem 1/4y+5=1/7y do I multiply everything by 7?